OpenStudy (anonymous):

out of pocket expenses for person in poor health were 3373 more than a person in good health. combined expenses for good and poor health were 6693. find out of pocket expenses for each person

OpenStudy (anonymous):

x+y = 6693 x=y+3373 2y + 3373 = 6693 2y = 3320 y = 1660 1660+3373 = 5033 poor health = $5033, good health = $1660